Read each problem carefully and clearly show all work. Indicate calculator functions used (ie T-Test, 2-SampTTest, etc) as well as calculator entries State your conclusions in terms of the context of the problem.

4) A researcher wants to compare the overall nutrient in children ages 2 to 19 that exclusively use skim milk instead of 1%, 2%, or whole milk. The researcher combined who consumed 1% or 2% milk into a “mixed milk” category. Is there significant evidence to support the belief that at least one mean nutrient intake of children is different from the others at the 10% significance level. The data and the Minitab output is below.

Skim Milk

Mixed Milk

Whole Milk

860

1006

880

850

1000

935

870

1020

845

900

1035

818

920

1024

870

886

1013

875

855

1060

881

860

1005

836

Analysis of Variance

Source

DF

Adj SS

Adj MS

F-Value

P-Value

Factor

2

118737

59368.6

78.48

<0.0001

Error

21

15887

756.5

Total

23

134624

         

a) Referring again to the problem above, if at least one mean is different, determine which type of milk drinker has a significantly different nutrient intake. Use the Tukey’s Test output below to explain your reasoning. Grouping Information Using the Tukey Method and 95% Confidence Factor N Mean Grouping Mixed 8 1020.38 A Skim 8 875.125 B Whole 8 867.50 B Means that do not share a letter are significantly different.
Answer #1

Test statistic:

F = 78.48

p-value = F.DIST.RT(78.4768, 2, 21) = 0.0000

Decision:

P-value < α, Reject the null hypothesis.

--

Tukey method:

Group Sample Mean Size
x̅1 875.125 8
x̅2 1020.375 8
x̅3 867.5 8

At α = 0.05, k = 3, N-K = 21, Q value = 3.56

Critical Range, CR = Q*√(MSW/n) = 3.56*√(756.5/8) = 34.62

If |difference| > CR, then the result is significant.

Comparison Diff. = (xi - xj) Critical Range Results
x̅1-x̅2 -145.25 34.62 Means are different
x̅1-x̅3 7.625 34.62 Means are not different
x̅2-x̅3 152.875 34.62 Means are different